Consider a charge of 10 nC at (1,0) cm, and a charge of -20 nC at (1,1) cm. Calculate the electric potential V at the origin. Remember that the potential contribution from each charge is a scalar. While the geometry of this problem is identical to a previous problem, that problem asked for the electric field at the origin, not the potential. This problem is considerably easier.
Answer #1

Q1 = 10nC , Q2= - 20nC

r1 =1cm = 0.01m, r2= (x^2+ y^2)^1/2=(2)^1/2cm = 0.0141m

V= kQ1/r1 + kQ2/r2

V= 9×10^9 ( 10/0.01 - 20/0.01414)× 10^-9= - 3729.84V